Given an array representation of a heap, with indexes starting at 0. If a child node is stored at index 50,
what index will contain its parent?
Answer #1
formula for index of parent is (child-1)/2
index of the parent is (50-1)/2 = 49/2 = 24

Answer: 24
index of the parent is 24